MGT 5 - MANAGERIAL ACCOUNTING PROBLEM 8 GIVEN: 66 VC/UNIT VC % FC TAX RATE AFTER-TAX NI 75.00% 396,000 30% 141,680 REQUIRED Pcan you help me understand the part that is required from the given?
Answer #1
% Per unit Totals Units = 27200
Sales 100% $                                        88 $ 23,93,600
VC 75% $                                        66 $ 17,95,200
CM 25% $                                        22 $   5,98,400 27200
FC $   3,96,000
PT 100% $   2,02,400
Tax 30% $       60,720
AT 70% $   1,41,680
Per unit
Sales = ($66 / 75%) X 100% = $                  88
CM = $88 - $66 = $                  22
Totals
PT = ($141680 / 70%) X 100% = $      2,02,400
Tax = ($141680 / 70%) X 30% = $          60,720
CM = $202400 + $396000 = $      5,98,400
Units = ($598400 / $22) = 27200
Sales = 27200 units X $88 = $    23,93,600
VC = 27200 units X $66 = $    17,95,200

Units = 27200 Sales Per unit =C4/B4 66 0.75 =B3-B4 =C3-C4 =D5/C5 5 CM 6 FC Totals =E5*C3 =E5*C4 =D7+D6 396000 =(D9/B9) *B7 =(